OpenStudy (anonymous):

Please help :) Factor the trinomial. x2 - 13x + 40

jimthompson5910 (jim_thompson5910):

What two numbers multiply to 40?

OpenStudy (anonymous):

20 times 2 or?

jimthompson5910 (jim_thompson5910):

20*2 = 40 is correct but notice how the factors 20 and 2 add to 20+2 = 22

jimthompson5910 (jim_thompson5910):

we want the factors of 40 to add to -13

jimthompson5910 (jim_thompson5910):

40 = 4*10 but 4+10 = 14 see how this works?

jimthompson5910 (jim_thompson5910):

you're looking for two numbers that multiply to 40 AND add to -13

OpenStudy (anonymous):

Yes :)I

jimthompson5910 (jim_thompson5910):

which two numbers do that

OpenStudy (tkhunny):

1*40 2*20 4*10 5*8 1+40 = 41 1-40 = -39 2+20 = 22 2-20 = -18 4+10 = 14 4-10 = -6 5+8 = 13 5-8 = 3 Did we get a 13 in there? All you have to do is hunt for it. No need for guessing.

OpenStudy (anonymous):

I don't get it

jimthompson5910 (jim_thompson5910):

notice how 8*5 = 40 and 8+5 = 13

jimthompson5910 (jim_thompson5910):

at the same time -8 * (-5) = 40 and -8 + (-5) = -13

jimthompson5910 (jim_thompson5910):

the fact that -8 * (-5) = 40 and -8 + (-5) = -13 means x^2 - 13x + 40 factors to (x-8)(x-5)

jimthompson5910 (jim_thompson5910):

notice how the -8 and -5 are in (x-8)(x-5)

OpenStudy (tkhunny):

That was every conceivable factorization of 40. That's all.